OpenStudy (anonymous):

solve the following equation to get p by itself: 4p-q=10

OpenStudy (anonymous):

p=(q+10)/4

OpenStudy (anonymous):

First add q to both sides of the equation, you get 4p = 10 + q Now, divide both sides of the equation by 4, you get p = (10 + q)/4